OpenStudy (anonymous):

put into standard form x+9=2(x-1)^2

OpenStudy (amistre64):

i that sposed to be "y+9" ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2x^2-5x-7=0

OpenStudy (amistre64):

well; we get: x+9 = 2(x^2 -2x +1) x+9 = 2x^2 -4x +4 0 = 2x^2 -5x -5

OpenStudy (amistre64):

if 2*1 = 4 that is lol

OpenStudy (anonymous):

2x^2-5x-7=0... so a=2,b=-5 and c=-7

OpenStudy (anonymous):

2x^2-5x-5=0 sorry for that its typing mistake